# Gravitational Waves Detected by a Burst Search in LIGO/Virgo’s Third Observing Run

### v1

@@ Abstract
Burst searches identify gravitational-wave (GW) signals in the detector data without use of a specific signal model, unlike the matched-filter searches that correlate data with simulated signal waveforms (templates). While matched filters are optimal for detection of known signals in the Gaussian noise, the burst searches can be more efficient in finding unusual events not covered by templates or those affected by non-Gaussian noise artifacts. Here, we report the detection of 3 gravitational wave signals that are uncovered by a burst search Coherent WaveBurst (cWB) optimized for the detection of binary black hole (BBH) mergers. They were found in the data from the LIGO/Virgo’s third observing run (O3) with a combined significance of 3.6 $\sigma$. Each event appears to be a BBH merger not previously reported by the LIGO/Virgo’s matched-filter searches. The most significant event has a reconstructed primary component in the upper mass gap ($m_{1}=70^{+36}_{-18}\,$), and unusually low mass ratio ($m_{2}/m_{1}\sim 0.3$), implying a dynamical or AGN origin. The 3 new events are consistent with the expected number of cWB-only detections in the O3 run ($4.8\pm 2.1$), and belong to the stellar-mass binary population with the total masses in the $70-100$ M range.
@@ 1 Introduction
Binary black hole (BBH) mergers can have a variety of origins that leave imprints on the binaries’ properties. Possible origins include isolated formation from stellar binaries (Dominik et al., 2012; Belczynski et al., 2016; de Mink & Mandel, 2016), dynamical formation in dense environments such as galactic nuclei or globular clusters (O’Leary et al., 2006; Banerjee et al., 2010; Rodriguez et al., 2016a), or gas capture in Active Galactic Nuclei (Bartos et al., 2017; Tagawa et al., 2020). Generally, isolated binaries are expected to be lower, close to equal masses with low spins aligned with the orbital angular momentum, while dynamical formation and gas-capture can produce heavier black holes through mass segregation and multiple consecutive mergers (Fishbach et al., 2017), more asymmetric masses due to the quasi-random pairing of the binary partners, and possibly high spins misaligned from the binary orbit (Rodriguez et al., 2016b; Talbot & Thrane, 2017). Dynamical formation and AGN gas-capture can also introduce orbital eccentricity as these binaries can interact soon before merger (Samsing et al., 2022).
Our ability to discover and identify black hole mergers depends mainly on the sensitivity of gravitational wave observatories, but also on the utilized search algorithms. Most binary searches use a gravitational-wave template bank that densely covers the targeted binary parameter space with waveforms, and use matched filters to compare each of the waveforms to the recorded data to find the best match (Usman et al., 2016; Messick et al., 2017; Aubin et al., 2021). Such template-based searches are optimal in the presence of stochastic noise, and if the covered parameter space includes that of the binary merger in question. Template-based gravitational wave searches have been exceedingly successful, and in fact, all gravitational wave signals discovered to date have been identified or co-identified by the template-based searches.
@@
To overcome the challenges of incomplete parameter space coverage and non-Gaussian detector noise, burst searches have been developed that do not closely rely on precise gravitational waveforms, but instead use more general assumptions about the signal and the background, e.g., by looking at excess energy coherently deposited in multiple gravitational wave detectors. A prominent model-agnostic search algorithm is coherent WaveBurst (cWB; Klimenko et al. (2008); Drago et al. (2020)). While cWB can detect binary mergers across the full parameter space accessible by the LIGO and Virgo detectors, its relative sensitivity compared to template-based methods makes it the most relevant for the heaviest black holes ($\gtrsim 20$ M; Abbott et al. (2022)), and in the case of unusual binary black holes not covered by template-based searches, such as eccentric and precessing mergers (Abac et al., 2023; Gadre et al., 2024).
@@ 2 Method
cWB is an analysis algorithm used in searches for transient GW signals with the networks of GW detectors. Designed to operate without a specific waveform model, cWB identifies coincident excess power in the multi-resolution time-frequency (TF) representations of the detector strain data (Klimenko et al., 2008, 2016), for signal frequencies up to a few kHz and duration up to a few seconds. After the excess power events are identified in the TF data, the cWB reconstructs the source sky location and signal waveforms recorded by the detectors with the constrained maximum likelihood method (Klimenko et al., 2005, 2016). The cWB detection statistic is based on the coherent energy $E_{c}$ obtained by cross-correlating the signal waveforms reconstructed in the detectors and normalized by the spectral amplitude of the noise. The $\sqrt{E_{c}}$ is the lower bound on the event network SNR $\rho_{N}$ and used for the initial selection of events produced by cWB. The production rate of cWB events is dominated by the non-stationary detector noise (glitches) and further reduced with the post-production vetoes. One of the primary signal-independent veto statistic is the network correlation coefficient $c_{c}=E_{c}/(E_{c}+E_{n})$, where $E_{n}$ is the normalized residual noise energy estimated after the reconstructed signal is subtracted from the data. Typically, for a GW signal $c_{c}\sim 1$, and for glitches $c_{c}<1$. Therefore, the candidate events with low value of $c_{c}$ are rejected as potential glitches. For GW signals the residual energy $E_{n}$ is expected to follow the $\chi^{2}$ distribution with the number of degrees of freedom $N_{DoF}$ which is proportional to the number of the TF data samples comprising the event. The reduced $\tilde{\chi}^{2}=E_{n}/N_{DoF}$ statistic is used as a powerful veto to identify and remove glitches with $\tilde{\chi}^{2}$ significantly larger than 1. While cWB uses the model-agnostic detection, reconstruction, and veto algorithms, it can also employ a weak signal-dependent veto to improve the identification of BBH signals. It is based on a generic property of a BBH signal which central frequency is $f_{c}\propto{1/M}$ and the slope of a chirp track in the TF domain is approximated as $\dot{f}\propto{{\cal M}^{5/3}_{\rm c}f^{11/3}}$ (Tiwari et al., 2015), where $M$ and ${\cal M}_{\rm c}$ are the red-shifted total and chirp masses respectively. Both parameters can be estimated from the frequency evolution $f(t)$ of detected events and used to improve the separation of BBH events and glitches.
@@
There are three significant changes in the cWB algorithm with respect to the version used during the O3 observing run (Abbott et al., 2023b).
@@
Finally, we substitute the post-production veto analysis with the machine-learning algorithm XGBoost. By using an ensemble of decision trees, it performs a classification of the event summary statistics produced by cWB ($E_{c},c_{c},\tilde{\chi},\rho_{e},\rho_{\times},r_{\times},r_{s},f_{c},{\cal M}_{\rm c}$, etc.) to improve separation of genuine GW events and glitches. The decision trees are trained on a representative set of simulated BBH events (section 2.2) and background events. Due to a weak dependence of the event summary statistics on the BBH dynamic, the algorithm retains a high detection efficiency for BBH signals outside of the training set. A detailed description of the XGBoost post-production analysis is published elsewhere (Mishra et al., 2021, 2022; Szczepańczyk et al., 2023). The final cWB detection statistic used for the estimation of the detection significance is the same as the one described in Mishra et al. (2022)
@@ 3 Results
In this section, we present the results from the reanalysis of the O3 data with the upgraded cWB pipeline. Unlike in the previous cWB run (Abbott et al., 2023b), which employed two different search configurations, the updated cWB performed a single search for both stellar-mass BBH and IMBH mergers in the detector frequency band of $16-320$ Hz. Employing a single search reduces the data processing time and eliminates the need for a trial factor. Due to the variation of the detector noise in the first (O3a) and the second (O3b) halves of the third observing run, the two XGBoost configurations were trained separately for the O3a and O3b data. The training procedure was similar to the method described in Mishra et al. (2022), with the updated list of the input summary statistics produced by cWB.
@@ 3.1 Detections
In Table 3 we report 33 GW events that are present in the GWTC-3 catalog (Abbott et al., 2023b), with an inverse false alarm rate (IFAR) of 1 year as the threshold for detection. In this search, while GW200219_094415 is missed (reported earlier with a low IFAR of $1.3$ yr), we detect 9 GW events that were previously missed by the O3 cWB search. In addition to the known GW events, cWB identified 3 new event candidates that were not previously reported in the GWTC-3 catalog. The GPS time and IFAR of the cWB-only event candidates, as well as their signal-to-noise ratios (SNR) and the probability of astrophysical origin ($P_{\mathrm{astro}}$) are shown in Table 1. The highest significance candidate has an IFAR $>40$ yr, placing it well above the detection threshold, while the lowest significance GW candidate, with IFAR $=4.5$ yr, is less clearly a detection. Nevertheless, the three new candidates together have a combined significance of 3.6 $\sigma$, representing a statistical excess of events detected only by the burst search.
@@ 3.2 Event properties
The parameter estimation results for the cWB candidates are reported in Table 2.
Our parameter estimation gives a good match for each of the candidates, as measured in the Bayes Factor (BF) of the best matching waveform for the events ($\ln(\mathrm{BF})>18$), indicating that a stellar-mass binary merger is an adequate explanation for each of the events. Fig 2 (left), shows the 1D & 2D posterior distribution for the component masses $m_{1}$ and $m_{2}$ of the source for each event with the IMRPhenomXPHM model. Fig 2 (right) shows the 1D & 2D posterior distribution for effective spin $\chi_{\mathrm{eff}}$ and effective precession spin $\chi_{\mathrm{p}}$ for each event with the IMRPhenomXPHM model. Based on these results, there is no evidence for precession. To confirm this finding, we also conducted parameter estimation using the aligned spin option of the IMRPhenomXPHM model. We observed that the BF ratio between the aligned and precessing spin results is approximately 1.

### v2

@@ Abstract
Burst searches identify gravitational-wave (GW) signals in the detector data without use of a specific signal model, unlike the matched-filter searches that correlate data with simulated signal waveforms (templates). While matched filters are optimal for detection of known signals in the Gaussian noise, the burst searches can be more efficient in finding unusual events not covered by templates or those affected by non-Gaussian noise artifacts. Here, we report the detection of 3 gravitational wave signals that are uncovered by a burst search Coherent WaveBurst (cWB) optimized for the detection of binary black hole (BBH) mergers. They were found in the data from the LIGO/Virgo’s third observing run (O3) with a combined significance of 3.6 $\sigma$. Each event appears to be a BBH merger not previously reported by the LIGO/Virgo’s matched-filter searches. The most significant event has a reconstructed primary component in the upper mass gap ($m_{1}=70^{+36}_{-18}\,$), and unusually low mass ratio ($m_{2}/m_{1}\sim 0.3$), implying a dynamical or AGN origin. The 3 new events are consistent with the expected number of cWB-only detections in the O3 run, and belong to the stellar-mass binary population with the total masses in the $70-100$ M range.
@@ 1 Introduction
Binary black hole (BBH) mergers can have a variety of origins that leave imprints on the binaries’ properties. Possible origins include isolated formation from stellar binaries (Dominik et al., 2012; Belczynski et al., 2016; de Mink & Mandel, 2016), binary evolution of Pop III stars (Tanikawa et al., 2021; Tanikawa, 2024; Mestichelli et al., 2024), dynamical formation in dense environments such as galactic nuclei or globular clusters (O’Leary et al., 2006; Banerjee et al., 2010; Rodriguez et al., 2016a), or gas capture in Active Galactic Nuclei (Bartos et al., 2017; Tagawa et al., 2020). Generally, isolated binaries are expected to be lighter, close to equal mass black holes with low spins aligned with the orbital angular momentum. While dynamical formation and gas-capture can produce heavier black holes through mass segregation and multiple consecutive mergers (Fishbach et al., 2017), more asymmetric masses due to the quasi-random pairing of the binary partners, and possibly high spins misaligned from the binary orbit (Rodriguez et al., 2016b; Talbot & Thrane, 2017). Dynamical formation and AGN gas-capture can also introduce orbital eccentricity as these binaries can interact soon before merger (Samsing et al., 2022).
Our ability to discover and identify black hole mergers depends mainly on the sensitivity of the GW detectors, but also on the utilized search algorithms. Most binary searches use a gravitational-wave template bank that densely covers the targeted binary parameter space with waveforms, and use matched filters to compare each of the waveforms to the recorded data to find the best match (Usman et al., 2016; Messick et al., 2017; Aubin et al., 2021). Such template-based searches are optimal in the presence of stochastic noise, and if the covered parameter space includes that of the binary merger in question. Template-based gravitational wave searches have been exceedingly successful, and in fact, all gravitational wave signals discovered to date have been identified or co-identified by the template-based searches.
@@
To overcome the challenges of incomplete parameter space coverage and non-Gaussian detector noise, burst searches have been developed that do not closely rely on precise gravitational waveforms, but instead use more general assumptions about the signal and the background, e.g., by looking at excess energy coherently deposited in multiple gravitational wave detectors. A prominent model-agnostic search algorithm is coherent WaveBurst (cWB; Klimenko et al. (2008, 2016)). While cWB can detect binary mergers across the full parameter space accessible by the LIGO and Virgo detectors, its relative sensitivity compared to template-based methods makes it the most relevant for the heaviest black holes ($\gtrsim 20$ M; Abbott et al. (2022)), and in the case of unusual binary black holes not covered by template-based searches, such as eccentric and precessing mergers (Abac et al., 2023; Gadre et al., 2024).
@@ 2 Method
cWB is an analysis algorithm used in searches for transient GW signals with the networks of GW detectors. Designed to operate without a specific waveform model, cWB identifies coincident excess power in the multi-resolution time-frequency (TF) representations of the detector strain data (Klimenko et al., 2008, 2016), for signal frequencies up to a few kHz and duration up to a few seconds. After the excess power events are identified in the TF data, the cWB reconstructs the source sky location and signal waveforms recorded by the detectors with the constrained maximum likelihood method (Klimenko et al., 2005, 2016). The cWB detection statistic is based on the coherent energy $E_{c}$ obtained by cross-correlating the signal waveforms reconstructed in the detectors and normalized by the spectral amplitude of the noise. The $\sqrt{E_{c}}$ is the lower bound on the event network SNR $\rho_{N}$ and used for the initial selection of events produced by cWB. The production rate of cWB events is dominated by the non-stationary detector noise (glitches) and further reduced with the post-production vetoes. One of the primary signal-independent veto statistic is the network correlation coefficient $c_{c}=E_{c}/(E_{c}+E_{n})$, where $E_{n}$ is the normalized residual noise energy estimated after the reconstructed signal is subtracted from the data. Typically, for a GW signal $c_{c}\sim 1$, and for glitches $c_{c}<1$. Therefore, the candidate events with low value of $c_{c}$ are rejected as potential glitches. For GW signals the residual energy $E_{n}$ is expected to follow the $\chi^{2}$ distribution with the number of degrees of freedom $N_{DoF}$ which is proportional to the number of the TF data samples comprising the event. The reduced $\tilde{\chi}^{2}=E_{n}/N_{DoF}$ statistic is used as a powerful veto to identify and remove glitches with $\tilde{\chi}^{2}$ significantly larger than 1. While cWB uses the model-agnostic detection, reconstruction, and veto algorithms, it can also employ a weak signal-dependent veto to improve the identification of BBH signals. It is based on a generic property of a BBH signal which merger frequency is $f_{c}\propto{1/M}$ and the slope of a chirp track in the TF domain is approximated as $\dot{f}\propto{{\cal M}^{5/3}_{\rm c}f^{11/3}}$ (Tiwari et al., 2015), where $M$ and ${\cal M}_{\rm c}$ are the red-shifted total and chirp masses respectively. Both parameters can be estimated from the frequency evolution $f(t)$ of detected events and used to improve the separation of BBH events and glitches.
@@
There are three significant changes in the cWB algorithm with respect to the version used during the O3 observing run (Abbott et al., 2023b; Drago et al., 2020).
@@
Finally, we substitute the post-production veto analysis with the machine-learning algorithm XGBoost. By using an ensemble of decision trees, it performs a classification of the event summary statistics produced by cWB ($E_{c},c_{c},\tilde{\chi},\rho_{e},\rho_{\times},r_{\times},r_{s},f_{c},{\cal M}_{\rm c}$, etc.) to improve separation of the genuine GW events and glitches. The decision trees are trained on a representative set of simulated BBH events (section 2.2) and background events. Due to a weak dependence of the event summary statistics on the BBH dynamic, the algorithm retains a high detection efficiency for BBH signals outside of the training set. A detailed description of the XGBoost post-production analysis is published elsewhere (Mishra et al., 2021, 2022; Szczepańczyk et al., 2023). The final cWB detection statistic used for the estimation of the detection significance is the same as the one described in Mishra et al. (2022)
@@ 3 Results
In this section, we present the results from the reanalysis of the O3 data with the upgraded cWB pipeline. Unlike in the previous cWB run (Abbott et al., 2023b), which employed two different search configurations, the updated cWB performed a single search for both stellar-mass BBH and IMBH mergers in the detector frequency band of $16-320$ Hz. Employing a single search reduces the data processing time and eliminates the need for a trial factor. Due to the variation of the detector noise in the first (O3a) and the second (O3b) halves of the O3 run, the two XGBoost configurations were trained separately for the O3a and O3b data. The training procedure was similar to the method described in Mishra et al. (2022), with the updated list of the input summary statistics produced by cWB.
@@ 3.1 Detections
In Table 3 we report 33 GW events that are present in the GWTC-3 catalog (Abbott et al., 2023b), with an inverse false alarm rate (IFAR) of 1 year as the threshold for detection. In this search we detect 9 GW events that were previously missed by the O3 cWB search. In addition to the known GW events, cWB identified 3 new event candidates that were not previously reported in the GWTC-3 catalog. The GPS time and IFAR of the cWB-only event candidates, as well as their signal-to-noise ratios (SNR) and the probability of astrophysical origin ($P_{\mathrm{astro}}$) are shown in Table 1. The highest significance candidate has an IFAR $>40$ yr, placing it well above the detection threshold, while the lowest significance GW candidate, with IFAR $=4.5$ yr, is less clearly a detection. Nevertheless, the three new candidates together have a combined significance of 3.6 $\sigma$, representing a statistical excess of events detected only by the burst search.
@@ 3.2 Event properties
The parameter estimation results for the cWB candidates are reported in Table 2. Our parameter estimation gives a good match for each of the candidates, as measured in the Bayes Factor (BF) of the best matching waveform for the events ($\ln(\mathrm{BF})>18$), indicating that a stellar-mass binary merger is an adequate explanation for each of the events. Fig 2 (left), shows the 1D & 2D posterior distribution for the component masses $m_{1}$ and $m_{2}$ of the source for each event with the IMRPhenomXPHM model. Fig 2 (right) shows the 1D & 2D posterior distribution for effective spin $\chi_{\mathrm{eff}}$ and effective precession spin $\chi_{\mathrm{p}}$ for each event with the IMRPhenomXPHM model. Based on these results, there is no evidence for precession. To confirm this finding, we also conducted parameter estimation using the aligned spin option of the IMRPhenomXPHM model. We observed that the BF ratio between the aligned and precessing spin results is approximately 1.